OpenStudy (anonymous):

find the derivative (x^4 + 3x^2 -2) ^5

OpenStudy (whpalmer4):

Do you know the chain rule?

OpenStudy (anonymous):

chain rule for this one \[\frac{d}{dx}[f^5(x)]=5f(x)\times f'(x)\]

OpenStudy (anonymous):

so far i have 5(x^4+3x^2-2)^4 * (4x^3+6x)

OpenStudy (anonymous):

i dont know how to keep going from there

OpenStudy (anonymous):

i would stop

OpenStudy (anonymous):

i need to fully simplify it completely

OpenStudy (whpalmer4):

That is fully simplified. If your problem said to expand it, well, the highest order term in that sucker after expansion is \(x^{19}\) (!)

OpenStudy (anonymous):

yeah there is no way in hell you want to multiply this out plus, there is no such mathematical operation as "simplify" this is a simple as it gets

OpenStudy (anonymous):

find the derivative (x^4 + 3x^2 -2) ^5 f'(x)=5(x^4+3x^2-2)^4(4x^3+6x)

OpenStudy (anonymous):

yeah i need to expand it.

OpenStudy (whpalmer4):

Okay, if you want to check your work, here's what it expands to: \[2560 x^3+3600 x^5+1520 x^7-4770 x^9-1140 x^{11}+2100 x^{13}+1280 x^{15}+270 x^{17}\] plus at least one term lopped off each end so no one complains that I gave you the answer to a ridiculous problem...

OpenStudy (anonymous):

thank you